OpenStudy (anonymous):

Add: (3x-1) + (1-3y) (x-y) (y-x)

OpenStudy (anonymous):

\[(3x-1)/(x-y) +(1-3y)/(y-x)=(3x-1)/(x-y)-(3y-1)/(x-y) =(3x-3y)/(x-y)=3\]

OpenStudy (anonymous):

\[(3x-1)/(x-y)-(3y-1)/(x-y)\]=(3x-3y)/(x-y) =3

OpenStudy (anonymous):

sorry the above math is wrong it should be (3x-1)/(x-y)+(3y-1)/(x-y)=(3x-3y-2)/(x-y)=3-[2/(x-y)]